Norachai and Kali are selling fruit for a school fundraiser. Customers can buy small boxes of tangerines and large boxes of tangerines. Norachai sold 5 small boxes of tangerines and 14 large boxs of tangerines for a total of $166. Kali sold 13 small boxes of tangerines and 7 large boxes of tangerines for a total of $167. Find the cost of each of one small box of tangerines and one large box of tangerines.
How would you solve that?
Thank you.

Answer by Cromlix(4381)

Small boxes = x
Large boxes = y
5x + 14y = 166...1
13x + 7y = 167...2
Multiply (2) by 2
5x + 14y = 166 ...1
26x + 14y = 334....3
Subtract (1) from (3)
21x = 168
x = 8
Substitute x = 8 into
equation (1)
5x + 14y = 166
5(8) + 14y = 166
40 + 14y = 166
14y = 126
y = 9.
Therefore a small box of tangerines
cost $8
A large box of tangerines
cost $9.
Hope this helps.
:-)